OpenStudy (anonymous):

6 - 9w = 24 please help

OpenStudy (ivettef365):

6 - 9 w = 24 -6 -6 =========== -9w = 18

OpenStudy (anonymous):

First collect your like terms. I would combine that 6 with the 24 somehow, so that your factor with a variable (-9w) is by itself.

OpenStudy (ivettef365):

then you have -9w =18 ---- --- -9 -9

OpenStudy (ivettef365):

can you do it from there?

OpenStudy (anonymous):

w=(-2)

OpenStudy (ivettef365):

Good job

OpenStudy (anonymous):

6-(9x(-2))=24

OpenStudy (anonymous):

its right isnt it @ivettef365

OpenStudy (anonymous):

yess